Flamo:

Find the distance between the points (13, 20) and (18, 8).

Flamo:

\[d = \sqrt{(18 - 13)^2 + (8 - 13^2)}\]

Flamo:

\[d = \sqrt{(5)^2 + (-5)^2}\]

Flamo:

\[d= \sqrt{25 + (-25)}\]

Ultrilliam:

@Shadow

Flamo:

Ty

Flamo:

\[d = \sqrt{25 + 25}\]

Flamo:

\[d = \sqrt{50}\]

Flamo:

7.07 = about 7.

Flamo:

Ye! I did it, xD By Myself..

Shadow:

Check y1

Flamo:

o

Flamo:

Ye, Its Correct, Right?

Shadow:

\[d = \sqrt (18 - 13)^2 + (8 - 20)^2\]

Shadow:

20 is y1, not 13

Shadow:

13 is x1

Flamo:

O

Flamo:

\[d = \sqrt{(18- 13)^2 + (8-20)^2} \]

Flamo:

\[d = \sqrt{(5)^2 + (-12)^2}\]

Flamo:

\[d = \sqrt{25 + (-144})\]

Flamo:

Right?

Shadow:

no

Flamo:

o

Shadow:

\[(-12)^2 = -12 \times -12 = 12 \times 12 = 144\]

Shadow:

A negative times a negative is a positive

Flamo:

Ye, I know that Because of the rule. But when Ever I check on a Calculator it says Negitive.

Flamo:

\[d = \sqrt{25 + 144}\]

Flamo:

\[d = \sqrt{169}\]

Flamo:

\[d = 13\]

Flamo:

Ye?